A body start from rest and accelerate uniformly at 5meter per seconds square. Calculate the time taken by the body to cover a distance of 1 km



1
Expert's answer
2021-03-29T08:59:53-0400

"S=at^2/2\ \to t=\sqrt{2S/a}=\sqrt{2\cdot 1000/5}=20\ (s)" . Answer